OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center (0, 8), radius = 8

jimthompson5910 (jim_thompson5910):

hint: (x-h)^2+(y-k)^2=r^2 is the general equation of a circle where (h,k) is the center r is the radius

OpenStudy (anonymous):

Im confused

jimthompson5910 (jim_thompson5910):

what is the center in this case?

OpenStudy (anonymous):

(0,8)

jimthompson5910 (jim_thompson5910):

so (h,k) = (0,8) which means h = 0 k = 8

jimthompson5910 (jim_thompson5910):

the radius is given to be 8, so r = 8

OpenStudy (anonymous):

so whats the answer

jimthompson5910 (jim_thompson5910):

plug in h = 0, k = 8 and r = 8 into (x-h)^2+(y-k)^2=r^2 to get your answer

OpenStudy (anonymous):

i got it thanks

jimthompson5910 (jim_thompson5910):

yw